OpenStudy (anonymous):

how many rounds are necessary to find a champion in a single elimination softball tournament that starts with 64 teams?

OpenStudy (mrnood):

there are 2 players in the final there are 4 players in the semi-final there are 8 players in the quarter keep counting till you get to 64

OpenStudy (anonymous):

I need to create an exponential function

OpenStudy (mrnood):

each round has half (0.5 times) the number of players as the previous..

OpenStudy (anonymous):

so 64(0.5)^x?

OpenStudy (mrnood):

yes

OpenStudy (anonymous):

Could you help me on one more question?

OpenStudy (anonymous):

@MrNood

OpenStudy (mrnood):

look above - the first round has 64 players 64 (0.5)^1 = 32 have another think....

OpenStudy (anonymous):

Five hours after drinking an energy shot, half of the initial caffeine is still present in the body. If an energy shot has 96mg of caffeine, how may hours till 12mg of it remaining?

OpenStudy (mrnood):

this is the same equation it's about half life - the factor is 0.5 again

OpenStudy (mrnood):

the solution is the number of 'half life' periods - so the number of hours is 5 times this

OpenStudy (anonymous):

96(0.5)^x 96(0.5)^3 =12 3*5=15?

OpenStudy (mrnood):

yes

OpenStudy (anonymous):

ok thanks, you've been a big help